OpenStudy (anonymous):

a contractor has a blueprint for a house drawn to the scale is 1in:3ft. A wall on the blueprint is 6.5 inches long. How long is the actual wall?

OpenStudy (anonymous):

1/3 = 6.5/x (1 in to 3 ft = 6.5 in to x ft) cross multiply (1)(x) = (3)(6.5) x = 19.5 ft check... 1/3 = 6.5/19.5 cross multiply (1)(19.5) = (3)(6.5) 19.5 = 19.5 (correct) 6.5 in = 19.5 ft (the measurement of the wall)